OpenStudy (anonymous):

A perfect square trinomial never consists of squared 1st term and third terms. True or false?

OpenStudy (anonymous):

a^2 + 2ab + b^2 = (a + b)^2,\,\! and a^2 - 2ab + b^2 = (a - b)^2.\,\!

OpenStudy (anonymous):

what do you think

OpenStudy (anonymous):

False?

OpenStudy (anonymous):

bingo